OpenStudy (andu1854):

What is a 3 -child family's probability of having two boys given that at least one of their children is a boy?

OpenStudy (mathstudent55):

In three births, these are the possibilities: BBB BBG BGB BGG GBB GBG GGB GGG

OpenStudy (andu1854):

so it would be 3/8?

OpenStudy (mathstudent55):

If you know that one child is a boy, then you can eliminate GGG

OpenStudy (mathstudent55):

There are 7 possibilities with at least one boy.

OpenStudy (mathstudent55):

There are 4 cases of at least two boys out of 7 cases with at least one boy.

OpenStudy (triciaal):

one approach is to do a "map" equal chance for boy or girl = 1/2 1/2*1/2*1/2 = 1/8 so there are 8 possible outcomes

OpenStudy (triciaal):